OpenStudy (anonymous):

fan and medal! The table below represents the displacement of a turtle from its nest as a function of time: Time (hours) x Displacement from nest (miles) y 0 5 1 27 2 49 3 71 4 93 Part C: What would be the domain of the function if the turtle continued to walk at this rate until it traveled 225 miles from the nest? (2 points)

OpenStudy (anonymous):

@welshfella @ganeshie8 @peachpi someone please help me

OpenStudy (nonopro):

hi

OpenStudy (nonopro):

oh

OpenStudy (anonymous):

@nonopro hey:( im still having trouble on this thing still

OpenStudy (anonymous):

do you know how to find the slope of a line?

OpenStudy (anonymous):

yea. the slope is: y1-y2/x1-x2

OpenStudy (anonymous):

right, so if you use 2 of your points, what slope will you get?

OpenStudy (anonymous):

ok... so my points are, 93 and 5 for the y. and for x, 4 and 0. will that work?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

93-5/4-0=88/4=22.

OpenStudy (anonymous):

right, and 5 is the y-intercept so the equation for the line is\[y=22x+5\]

OpenStudy (anonymous):

So now we plug in 225 for y and solve for x. \[225=22x+5\]

OpenStudy (anonymous):

so... what would I put for the x?

OpenStudy (anonymous):

don't put anything for x. We have to solve the equation to find it.

OpenStudy (anonymous):

i agree with peachipi

OpenStudy (anonymous):

x is last

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

and whatever you do to one side you have to do to the other side

OpenStudy (anonymous):

so what should I do now?

OpenStudy (anonymous):

Subtract 5 from both sides